If 10.0 L of oxygen at STP are heated to 512 degrees C what will be the new volume of gas if the pressure is also increased to 202.6kPa?
1
Expert's answer
2020-04-18T07:06:47-0400

Solution:

The Ideal Gas Law can be written as: PV = nRT,

where P, V and T are the pressure, volume and temperature; n is the amount of substance; and R is the ideal gas constant.

Assuming 2 set of conditions:

Initial case: PiVi = niRTi

Final case: PfVf = nfRTf

Since the ideal gas constant (R) has the same value in each case, one will get the final equation and the General Gas Equation:

(PiVi / niTi) = (PfVf / nfTf).

In our case, the amount of substance (n) does not change, therefore:

(PiVi / Ti) = (PfVf / Tf),

or PiViTf = PfVfTi


Pi = 101.325 kPa (1atm, STP); Vi = 10.0 L; Ti = 273.15 K (0oC, STP);

Pf = 202.6 kPa; Vf = unknown; Tf = 512oC = 785.15 K.


PiViTf = PfVfTi

Then,

(101.325 kPa * 10.0 L * 785.15 K) = (202.6 kPa * Vf * 273.15 K).

Vf = (101.325 kPa * 10.0 L * 785.15 K) / (202.6 kPa * 273.15 K) = 14.37 L

Vf = 14.37 L = 14.4 L


Answer: 14.4 L is the new volume (Vf) of gas.